OpenStudy (anonymous):

Given that f(x)=(x-2)(x^3+ax^2+bx+c). Find values of a,b,c.

OpenStudy (anonymous):

Is that the entire problem?

OpenStudy (anonymous):

oh, i just realised its given that f(x)=x^4-4x-4x-8

OpenStudy (anonymous):

so would you do that one divide by x-2?

OpenStudy (campbell_st):

well you need to do polynomial or synthetic division to get the coefficients a, b and c

OpenStudy (campbell_st):

opps the original function is missing some info... powers of x

OpenStudy (anonymous):

no, thats what the question says. I guess i would have to use 0x^3 and 0x^2 to fill in the blanks to divide

OpenStudy (campbell_st):

you have \[x^4 - 4x - 4x -8 = x^4 - 8x - 8\] so I think there is a power missing

OpenStudy (anonymous):

i dunno, thats what it says on the paper?